OpenStudy (anonymous):

Solve using elimination: 7x − 3y − 1= 0 4y − 6x = 12 A. (6, 2) B. (7,18) C. (4, 9) D. (0, -1/3)

OpenStudy (anonymous):

Answer is C...

OpenStudy (maheshmeghwal9):

@waterineyes do not give askers the answers only. make them understand it:)

OpenStudy (anonymous):

You might want to explain how you got that so Senior can understand and maybe try to solve one.

OpenStudy (anonymous):

He's had multiple equations just like this explained to him. Asked Senior012 to post how he thinks it should be solved, but no response yet.

OpenStudy (campbell_st):

rewrite both equations 7x - 3y = 1 (1) -6x +4y = 12 (2) then equation (1) times 6 plus equation (2) times 7 to eliminate x 42x - 18y = 6 -42x +28y = 84 ---------------- 10x = 90 x = 9 substitute into either equation to find y

OpenStudy (campbell_st):

oops y = 9.... substitute into either equation to find x